Any ideas. I'm stuck and need help
A 20Ohm coil of 15 turns and radius 10.0 cm surrou
Answer #1

Ans :- here R = 10cm N= 15 turns ; r = 5cm; n= 1000; I= 6 sin(120t)

Because the solenoid is long, we can pretend it is infinite, so all the magnetic field is contained inside the solenoid, and there is no magnetic field outside

The field inside the solenoid is given by

B = µ0nI ...........................1

So the flux through the large coil is

ΦB = ᶴ BdA = Nπr^2B = Nπr^2 µ0nI........................2

The induced emf is then

ε = − dΦB/ dt = −πµ0nNr2 dI /dt =- 3.14*4*3.14*10^-7*1000*15*(0.05)^2*6*120*cos(120t)

ε =106.48cos(120t)mV

t=3s

ε= 106.48mV